OpenStudy (anonymous):

in a survey of 100 students, the number of students studying the various languages were found to be as follows: English only 18, English but not Hindi 23, English and French 8, English 26, French 48, French and Hindi 8 and No language 24. How many students were studying Hindi and How many students were studying English and Hindi?

OpenStudy (anonymous):

Is it multiple choice?

OpenStudy (kirbykirby):

Let \(E\) = English \(H\) = Hindi \(F \) = French \(n(E \cap \overline{F} \cap \overline{H})=18\) \(n(E \cap \overline{H})=23\) \(n(E \cap F)= 8\) \(n(E)=26\) \(n(F)=48\) \(n(F\cap H)=8\) \(n(\overline{E} \cap \overline{F} \cap\overline{H} )=24\) \(n(E \cap H) = ?\) So: \(n(\overline{E} \cap \overline{F} \cap\overline{H} )=n(\overline{E \cup F \cup H})\) by DeMorgan's Law. \(n(\overline{E \cup F \cup H})=1-n(E \cup F \cup H)\) by complements. \[ n(E \cup F \cup H)=n(E)+n(F)+n(H)-n(E \cap F)-n(E \cap H) \\~~~~~~~~~~~~~~~~~~~~~~~~~~-n(F \cap H) + n(E \cap F \cap H)\\ ~ \\100-24 =1-[26+48+n(H)-8-n(E \cap H)-8+n(E \cap F \cap H)]\] Notice though that \(n(E)=n(E \cap H)+ n(E \cap \overline {H})\)  giving \(n(E \cap H)=3+0=3\)

OpenStudy (anonymous):

Woah.. I am so sorry// I don't understand this at all .

OpenStudy (kirbykirby):

oops that should say 100 - 76 on the before last line